# Agnostic Process Tomography

### v1

@@ 4.1 Bounded-gate circuits
In this section, we consider AUE when the concept class $\mathcal{C}$ consists of unitaries with a bounded gate complexity. In particular, we consider $n$-qubit unitaries generated by $G$ two-qubit gates. This class of unitaries was studied in the realizable case in [27].
First, we prove a more general result about AUE for any class $\mathcal{C}$ of unitaries with finite size $|\mathcal{C}|<+\infty$. In the following proposition, we show that AUE with respect to the average-case distance metric $\mathrm{d_{avg}}$ and the class $\mathcal{C}$ can be accomplished using a sample complexity scaling as $\log|\mathcal{C}|$. Thus, if $\mathcal{C}$ is any class of unitaries generated by $\mathrm{poly}(n)$ gates, $\log|\mathcal{C}|\leq\mathrm{poly}(n)$ so that the sample complexity is efficient. Qualitatively, this says that efficient AUE is easy from a sample complexity perspective, but it can be difficult to achieve with respect to time complexity. This is similar in spirit to an analogous argument by [42] for agnostic state tomography. Their argument relies on shadow tomography [94], while ours is based on classical shadows and takes inspiration from a proof of [27].
@@
We note that the time complexity in the proposition depends on the complexity of computing the inner products of pairs of states. This is due to the use of Clifford classical shadows [111], which is not computationally efficient for arbitrary observables (see Lemma 13). The resulting complexity depends on the class of unitaries one aims to learn and how efficiently their output states can be represented. However, since the time complexity already scales linearly with the size of the concept class, this algorithm would be computationally inefficient for most classes of interest, so we do not analyze the time complexity in detail.
@@
The algorithm is the same as that of Section C.2(a) of [27]. Namely, we use Clifford classical shadows [111] to estimate the distance $\mathrm{d_{avg}}(U_{i},U)$ between the unknown unitary $U$ and all unitaries $U_{i}\in\mathcal{C}$. Then, our algorithm outputs the one with the minimal distance.
The analysis is the same as Section C.2(a) of [27], so we do not reproduce the full argument here. We start by preparing random input states from the ensemble
[…]
of random tensor products of single-qubit stabilizer states. Then, we use classical shadows [111] to estimate distances with respect to this ensemble. Now, by Equation (C.8) of [27], using
@@
$|d_{P}(U_{i},U)^{2}-(1-\hat{o}_{i})|\leq\epsilon^{\prime},\quad\forall i\in[|\mathcal{C}|]$ (4.5)
with probability at least $1-\delta$ for some $\delta,\epsilon^{\prime}>0$. Then, we can select $i^{\star}=\arg\min_{i}(1-\hat{o}_{i})$ and output $U_{i^{\star}}$, i.e., the $U_{i}\in\mathcal{C}$ with the smallest estimated distance from $U$ with respect to $d_{P}$. To relate this back to $\mathrm{d_{avg}}$, recall that $P$ forms a $3$-design [137, 138, 139, 140, 141], implying that the average infidelity over $P$ is the same as that over the Haar measure. Thus for any $2^{n}$-dimensional unitaries $U_{1},U_{2}$,
@@
$W^{*}\triangleq\operatorname*{arg\,min}_{W\in\mathcal{C}}\mathrm{d_{avg}}(U,W)\quad\text{ and }\quad\mathsf{opt}(U,\mathcal{C})=\mathrm{d_{avg}}(U,W^{*}).$ (4.23)
@@
Finally, we can obtain a sample complexity upper bound for AUE of unitaries with bounded gate complexity. This follows immediately from Lemma 9, which gives a bound on the metric entropy of unitaries with bounded gate complexity. However, by Corollary 1, the time complexity scales with the cardinality of the covering net, which is exponentially large (Lemma 9). This is not surprising, as even in the realizable case, there are lower bounds on the computational complexity of learning this class of unitaries [27] based on the hardness of RingLWE [142].
@@ 4.2 AUE from agnostic state tomography
Definition 14 (Agnostic state tomography [42]).
Let $10<\epsilon,\delta<$ 1. Let $\mathcal{D}$ be a class of quantum states. A learning algorithm $\mathcal{A}$ is an $(1,\epsilon,\delta)$-agnostic learner with respect to fidelity and the concept class $\mathcal{D}$ if, given access to copies of an arbitrary quantum state $\rho$, $\mathcal{A}$ outputs a hypothesis state $\rho^{\prime}$ such that
@@
Algorithms for agnostic state tomography with respect to fidelity are known for a variety of classes, e.g., stabilizer product states [42], stabilizer states, states with high stabilizer dimension, and discrete product states [43].
@@
Suppose there exists a learning algorithm for (proper) $(1,\epsilon^{2},\delta)$-agnostic state tomography with respect to fidelity $F$ for some class $\mathcal{D}$ of pure states. Let $\mathcal{C}$ be a concept class of unitaries. If |$\ket{v(U)}\in\mathcal{D}$ for all $U\in\mathcal{C}$, then there exists a learning algorithm for improper $(1,\epsilon,\delta)$-agnostic unitary estimation with respect to error in normalized Frobenius distance $d_{F}$ for the class $\mathcal{C}$. Moreover, this requires the same sample and time complexity as the $(1,\epsilon^{2},\delta)$-agnostic state tomography algorithm.
We note that our result holds for agnostically learning the unitary channel associated to the unknown unitary. Namely, if $U$ is our unknown unitary, then our algorithm agnostically learns the channel $U(\cdot)U^{\dagger}$ with respect to $d_{F}$ (the normalized Frobenius distance for superoperators from Definition 10). We note that this does not take into account any phase differences and so is slightly weaker than learning the unitary $U$ itself with respect to $d_{F}$ (the normalized Frobenius distance for operators from Definition 9). This is because our algorithm simply applies agnostic state tomography to the Choi state of the unknown process, inspired by [87], which does not contain this phase information.
Moreover, our analysis can be seen as extending that of [87] to the agnostic setting with respect to the normalized Frobenius distance $d_{F}$, rather than the (trivial) metric of entanglement infidelity defined as the infidelity between two Choi states:
@@
where we adopt the (non-standard) notation of [6] and $v(\Phi)$ denotes the Choi state of $\Phi$. We remark that we can also trivially obtain $(1,\epsilon,\delta)$-agnostic unitary estimation with respect to the entanglement infidelity from $(1,\epsilon,\delta)$-agnostic state tomography with respect to fidelity (notably, without the quadratic blowup with respect to $\epsilon$) under the same compatibility condition. Entanglement infidelity has been considered widely in (realizable) unitary estimation [87, 6, 88, 89, 90, 91, 92, 93]. However, we focus on the normalized Frobenius distance $d_{F}$ for consistency with our other results.
@@
Let $U$ denote the unknown unitary. The algorithm is simple: first, prepare the Choi state |$\ket{v(U)}$ of $U$ and then apply the $(1,\epsilon^{2},\delta)$-agnostic state tomography algorithm to |$\ket{v(U)}$. The output of our algorithm is the density matrix of the pure state output by the state tomography algorithm.
The agnostic state tomography algorithm outputs some state |$\ket{\psi}$ such that
$F(v(U),\outerproduct{\psi}{\psi})\geq\max_{\ket{\phi}\in\mathcal{D}}F(v(U),\outerproduct{\phi}{\phi})-\epsilon^{2},$ (4.31)
where we use $v(U)\triangleq\outerproduct{v(U)}{v(U)}$ | to denote the density matrix corresponding to the pure Choi state. Note that this is consistent with the Choi state defined by the channel $\mathcal{E}(\rho)=U\rho U^{\dagger}$. We can equivalently write the above guarantee in terms of the infidelity
$1-\left|\innerproduct{v(U)}{\psi}\right|^{2}\leq\min_{\ket{\phi}\in\mathcal{D}}\left(1-\left|\innerproduct{v(U)}{\phi}\right|^{2}\right)+\epsilon^{2}.$ (4.32)
@@
by definition of $d_{F}$ (Definition 9), where |$\ket{\phi},\ket{\psi}$ are each on $2n$ qubits. Then, the above guarantee can be written as
$4^{n}d_{F}(v(U),\outerproduct{\psi}{\psi})^{2}\leq 4^{n}\min_{\ket{\phi}\in\mathcal{D}}d_{F}(v(U),\outerproduct{\phi}{\phi})^{2}+\epsilon^{2}.$ (4.35)
@@
$2^{n}d_{F}(v(U),\outerproduct{\psi}{\psi})\leq 2^{n}\min_{\ket{\phi}\in\mathcal{D}}d_{F}(v(U),\outerproduct{\phi}{\phi})+\epsilon.$ (4.36)
Moreover, because |$\ket{v(V)}\in\mathcal{D}$ for all $V\in\mathcal{C}$, we can upper bound this minimum by the minimum over all $V$ in $\mathcal{C}$ instead. Thus, we have
$2^{n}d_{F}(v(U),\outerproduct{\psi}{\psi})\leq 2^{n}\min_{V\in\mathcal{C}}d_{F}(v(U),v(V))+\epsilon.$ (4.37)
@@
$d_{F}(\mathcal{U}_{1},\mathcal{U}_{2})$ $=\frac{1}{2^{n}\sqrt{2}}\norm{\mathcal{U}_{1}-\mathcal{U}_{2}}_{F}$ (4.38)
$=\frac{1}{2^{n}\sqrt{2}}\norm{J(U_{1})-J(U_{2})}_{F}$ (4.39)
$=\frac{1}{\sqrt{2}}\norm{v(U_{1})-v(U_{2})}_{F}$ (4.40)
@@
Here, we use $\mathcal{U}$ to denote the unitary channel $U(\cdot)U^{\dagger}$. In the first and second lines, we use the definition of $d_{F}$ for superoperators (Definition 10). In the third line, we use that the Choi representation is simply the unnormalized Choi state. Here, recall that $v(U)=\outerproduct{v(U)}{v(U)}$ |, so the normalization factor is $2^{n}$ for this density matrix. Finally, in the last line, we use the definition of the normalized Frobenius distance for operators (Definition 9). Thus, this tells us that, indeed, agnostically learning the Choi state of a unitary with respect to $2^{n}d_{F}$ is the same as learning the corresponding unitary channel with respect to $d_{F}$. Hence, Equation 4.37 is exactly the guarantee needed for improper $(1,\epsilon,\delta)$-AUE with respect to the class $\mathcal{C}$. ∎
This gives a fairly general result for converting an agnostic state tomography algorithm with respect to $\mathcal{D}$ into an improper agnostic unitary estimation algorithm with respect to $\mathcal{C}$ when |$\ket{v(U)}\in\mathcal{D}$ for all $U\in\mathcal{C}$. As remarked earlier, this holds for many classes of unitaries for which there already exist agnostic state tomography algorithms with respect to a suitable class. In particular, there exist agnostic state tomography algorithms for stabilizer states, states with high stabilizer dimension, and discrete product states [42, 43]. Note that the corresponding classes of unitaries, i.e., Clifford circuits [143], Clifford + T circuits, and unitaries consisting of a tensor product of single-qubit gates, respectively, all satisfy the required property of Proposition 2. Thus, we can directly apply Proposition 2 with the state-of-the-art sample/time complexity bounds for agnostic state tomography to obtain improper AUE algorithms for the respective unitary classes. These are each direct corollaries of Corollary 6.3, Theorem 7.1, Corollary 8.3, and Corollary 9.3 in [43].
Corollary 3 (Improper AUE of Clifford Circuits; via Corollary 6.3 in [43]).
@@
Corollary 4 (Improper AUE of Clifford + T Circuits; via Theorem 7.1 in [43]).
@@
Corollary 5 (Improper AUE of Product Circuits; via Corollary 8.3 in [43]).
Let $\mathcal{K}$ be a set of single-qubit gates for which $|\innerproduct{v(U_{1})}{v(U_{2})}|^{2}\leq 1-\mu$ for any distinct $U_{1},U_{2}\in\mathcal{K}$. There exists a learning algorithm for improper $(1,\epsilon,\delta)$-agnostic unitary estimation of the class $\mathcal{C}=\mathcal{K}^{\otimes n}$ with respect to error in $d_{F}$ using
@@
Corollary 6 (Improper AUE of Clifford Product Circuits; via Corollary 9.3 in [43]).
@@ 4.3 Pauli strings
1:Prepare $N$ copies of $|v(U)\rangle$ and measure them in the $\{|v(\sigma_{x})\rangle\}_{x}$ basis to get strings $\{x_{i}\}_{i=1}^{N}$
@@
$\left|\hat{p}(x)-|\hat{U}_{x}|^{2}\right|\leq\epsilon_{1},\quad\forall x\in\mathcal{P}_{n}$ (4.59)
@@
$|\hat{U}_{z}|^{2}+\epsilon_{1}\geq\hat{p}(z)\quad\text{ and }\quad\hat{p}(x^{\star})\geq|\hat{U}_{x^{\star}}|^{2}-\epsilon_{1}.$ (4.61)
@@ 4.4 Unitary $k$-juntas
Now, we consider agnostic learning with respect to the class of $n$-qubit unitary $k$-juntas $\mathcal{J}_{k,n}$ [28]. An $n$-qubit $k$-junta is a $2^{n}$-dimensional unitary that acts nontrivially on $k$ qubits. More formally, we have the following definition.
@@
The algorithm is reminiscent of other junta learning algorithms [24, 144] in structure. Namely, we first identify the qubits $S$ that the $k$-junta acts nontrivially on and then learn the unitary on the set $S$. In contrast to [24], we use the Pauli weight (Definition 6) rather than the influence to measure how nontrivially a unitary acts on a given qubit. In the realizable setting, both the weight and influence of a subset are equivalent. However, in the agnostic setting, these notions are quite distinct, as demonstrated by the following example.
@@
$\mathcal{P}_{S}=\{x\in\{0,1,2,3\}^{n}:\mathsf{supp}(x)\subseteq S\}$ (4.68)
for a subset $S\subseteq[n]$. The resulting learned operator $V_{1}=\sum_{x\in\mathcal{P}_{S}}\alpha_{x}\sigma_{x}$ may not be unitary, so in order to achieve proper agnostic unitary estimation, we need an additional projection which approximates $V_{1}$ by a unitary. For this projection, we use Lemma 14, which takes influence from [23]. This algorithm achieves the following guarantee.
@@
$|\hat{w}_{S_{j}}-w_{S_{j}}(U)|\leq\epsilon\quad\forall j\in[M]$ (4.71)
@@
$\hat{w}_{S}\triangleq\frac{1}{N}\sum_{i\in[N]}b_{S,i},\quad b_{S,i}\triangleq\mathbf{1}[x_{i}\in\mathcal{P}_{S}].$ (4.72)
@@
Here, the third line follows because measuring |$\ket{v(U)}$ in the basis {|$\{\ket{v(\sigma_{x})}\}_{x\in\mathcal{P}_{n}}$ gives outcome $x$ with probability $|\hat{U}_{x}|^{2}$. Thus, using
@@
$|\hat{w}_{S_{j}}-w_{S_{j}}(U)|\leq\epsilon\quad\forall j\in[M]$ (4.78)
@@
with probability at least $21-\delta/$ 2, for all subsets $S\subseteq[n]$ of cardinality $|S|=k$. By Lemma 15, this requires
@@
In the next step of Algorithm 2, we estimate all $4^{k}$ Pauli coefficients for all Pauli strings supported on $S^{\prime}$. We use Lemma 11 $4^{k}$ times, to obtain estimates $\alpha_{x}$ such that with probability at least $21-\delta/$ 2,
$|\alpha_{x}-\hat{U}_{x}|\leq\epsilon_{2},\quad\forall x\in\mathcal{P}_{S^{\prime}}$ (4.91)
@@
$w_{S^{\prime}}(U)+\epsilon_{1}\geq\hat{w}_{S^{\prime}}\quad\text{ and }\quad\hat{w}_{S^{\star}}\geq w_{S^{\star}}(U)-\epsilon_{1}.$ (4.106)

### v2

@@ 4.1 Bounded-gate circuits
In this section, we consider AUE when the concept class $\mathcal{C}$ consists of unitaries with a bounded gate complexity. In particular, we consider $n$-qubit unitaries generated by $G$ two-qubit gates. This class of unitaries was studied in the realizable case in [zhao2023learning].
First, we prove a more general result about AUE for any class $\mathcal{C}$ of unitaries with finite size $|\mathcal{C}|<+\infty$. In the following proposition, we show that AUE with respect to the average-case distance metric $\mathrm{d_{avg}}$ and the class $\mathcal{C}$ can be accomplished using a sample complexity scaling as $\log|\mathcal{C}|$. Thus, if $\mathcal{C}$ is any class of unitaries generated by $\mathrm{poly}(n)$ gates, $\log|\mathcal{C}|\leq\mathrm{poly}(n)$ so that the sample complexity is efficient. Qualitatively, this says that efficient AUE is easy from a sample complexity perspective, but it can be difficult to achieve with respect to time complexity. This is similar in spirit to an analogous argument by [grewal2024agnostic] for agnostic state tomography. Their argument relies on shadow tomography [aaronson2018shadow], while ours is based on classical shadows and takes inspiration from a proof of [zhao2023learning].
@@
We note that the time complexity in the proposition depends on the complexity of computing the inner products of pairs of states. This is due to the use of Clifford classical shadows [huang2020predicting], which is not computationally efficient for arbitrary observables (see Lemma 13). The resulting complexity depends on the class of unitaries one aims to learn and how efficiently their output states can be represented. However, since the time complexity already scales linearly with the size of the concept class, this algorithm would be computationally inefficient for most classes of interest, so we do not analyze the time complexity in detail.
@@
The algorithm is the same as that of Section C.2(a) of [zhao2023learning]. Namely, we use Clifford classical shadows [huang2020predicting] to estimate the distance $\mathrm{d_{avg}}(U_{i},U)$ between the unknown unitary $U$ and all unitaries $U_{i}\in\mathcal{C}$. Then, our algorithm outputs the one with the minimal distance.
The analysis is the same as Section C.2(a) of [zhao2023learning], so we do not reproduce the full argument here. We start by preparing random input states from the ensemble
$P=\mathrm{Uniform}(\{|\psi\rangle=U|0\rangle,U\in\mathrm{Cl}(n)\})$ (4.3)
of random stabilizer states, where $\mathrm{Cl}(n)$ denotes the $n$-qubit Clifford group. Then, we use classical shadows [huang2020predicting] to estimate distances with respect to this ensemble.
While [zhao2023learning] used random stabilizer product states as input, we use random stabilizer states instead. We note that changing the input distribution only changes the ensemble over which the distance is estimated, not the sample complexity for successful estimation. Now, by Equation (C.8) of [zhao2023learning], using
@@
$|d_{P}(U_{i},U)^{2}-(1-\hat{o}_{i})|\leq\epsilon^{\prime},\qquad\forall i\in[|\mathcal{C}|]$ (4.5)
with probability at least $1-\delta$ for some $\delta,\epsilon^{\prime}>0$. Then, we can select $i^{\star}=\arg\min_{i}(1-\hat{o}_{i})$ and output $U_{i^{\star}}$, i.e., the $U_{i}\in\mathcal{C}$ with the smallest estimated distance from $U$ with respect to $d_{P}$. To relate this back to $\mathrm{d_{avg}}$, recall that $P$ forms a $3$-design [kueng2015qubit, webb2015clifford, zhu2017multiqubit], implying that the average infidelity over $P$ is the same as that over the Haar measure. Thus for any $2^{n}$-dimensional unitaries $U_{1},U_{2}$,
@@
$W^{*}\triangleq\operatorname*{arg\,min}_{W\in\mathcal{C}}\mathrm{d_{avg}}(U,W)\qquad\text{ and }\qquad\mathsf{opt}(U,\mathcal{C})=\mathrm{d_{avg}}(U,W^{*}).$ (4.23)
@@
Finally, we can obtain a sample complexity upper bound for AUE of unitaries with bounded gate complexity. This follows immediately from Lemma 9, which gives a bound on the metric entropy of unitaries with bounded gate complexity. However, by Corollary 1, the time complexity scales with the cardinality of the covering net, which is exponentially large (Lemma 9). This is not surprising, as even in the realizable case, there are lower bounds on the computational complexity of learning this class of unitaries [zhao2023learning] based on the hardness of RingLWE [lyubashevsky2010ideal].
@@ 4.2 AUE from agnostic state tomography
Definition 14 (Agnostic state tomography [grewal2024agnostic]).
Let $0<\epsilon,\delta<1$. Let $\mathcal{D}$ be a class of quantum states. A learning algorithm $\mathcal{A}$ is an $(1,\epsilon,\delta)$-agnostic learner with respect to fidelity and the concept class $\mathcal{D}$ if, given access to copies of an arbitrary quantum state $\rho$, $\mathcal{A}$ outputs a hypothesis state $\rho^{\prime}$ such that
@@
Algorithms for agnostic state tomography with respect to fidelity are known for a variety of classes, e.g., stabilizer product states [grewal2024agnostic], stabilizer states, states with high stabilizer dimension, and discrete product states [chen2024stabilizer].
@@
[…]
We note that our result holds for agnostically learning the unitary channel associated to the unknown unitary. Namely, if $U$ is our unknown unitary, then our algorithm agnostically learns the channel $U(\cdot)U^{\dagger}$ with respect to $d_{F}$ (the normalized Frobenius distance for superoperators from Definition 10). We note that this does not take into account any phase differences and so is slightly weaker than learning the unitary $U$ itself with respect to $d_{F}$ (the normalized Frobenius distance for operators from Definition 9). This is because our algorithm simply applies agnostic state tomography to the Choi state of the unknown process, inspired by [leung2000towards], which does not contain this phase information.
Moreover, our analysis can be seen as extending that of [leung2000towards] to the agnostic setting with respect to the normalized Frobenius distance $d_{F}$, rather than the (trivial) metric of entanglement infidelity defined as the infidelity between two Choi states:
@@
where we adopt the (non-standard) notation of [haah2023query] and $v(\Phi)$ denotes the Choi state of $\Phi$. We remark that we can also trivially obtain $(1,\epsilon,\delta)$-agnostic unitary estimation with respect to the entanglement infidelity from $(1,\epsilon,\delta)$-agnostic state tomography with respect to fidelity (notably, without the quadratic blowup with respect to $\epsilon$) under the same compatibility condition. Entanglement infidelity has been considered widely in (realizable) unitary estimation [leung2000towards, haah2023query, acin2001optimal, peres2002covariant, hayashi2006parallel, chiribella2005optimal, kahn2007fast, yang2020optimal]. However, we focus on the normalized Frobenius distance $d_{F}$ for consistency with our other results.
@@
[…]
where we use $|v(U)\triangleq\outerproduct{v(U)}{v(U)}$ to denote the density matrix corresponding to the pure Choi state. Note that this is consistent with the Choi state defined by the channel $\mathcal{E}(\rho)=U\rho U^{\dagger}$. We can equivalently write the above guarantee in terms of the infidelity
[…]
@@
[…]
@@
[…]
@@
$d_{F}(\mathcal{U}_{1},\mathcal{U}_{2})$ $=\frac{1}{2^{n}\sqrt{2}}\mathinner{\!\left\lVert\mathcal{U}_{1}-\mathcal{U}_{2}\right\rVert}_{F}$ (4.38)
$=\frac{1}{2^{n}\sqrt{2}}\mathinner{\!\left\lVert J(U_{1})-J(U_{2})\right\rVert}_{F}$ (4.39)
$=\frac{1}{\sqrt{2}}\mathinner{\!\left\lVert v(U_{1})-v(U_{2})\right\rVert}_{F}$ (4.40)
@@
Here, we use $\mathcal{U}$ to denote the unitary channel $U(\cdot)U^{\dagger}$. In the first and second lines, we use the definition of $d_{F}$ for superoperators (Definition 10). In the third line, we use that the Choi representation is simply the unnormalized Choi state. Here, recall that $|v(U)=\outerproduct{v(U)}{v(U)}$, so the normalization factor is $2^{n}$ for this density matrix. Finally, in the last line, we use the definition of the normalized Frobenius distance for operators (Definition 9). Thus, this tells us that, indeed, agnostically learning the Choi state of a unitary with respect to $2^{n}d_{F}$ is the same as learning the corresponding unitary channel with respect to $d_{F}$. Hence, Equation 4.37 is exactly the guarantee needed for improper $(1,\epsilon,\delta)$-AUE with respect to the class $\mathcal{C}$. ∎
[…]
Corollary 3 (Improper AUE of Clifford Circuits; via Corollary 6.3 in [chen2024stabilizer]).
@@
Corollary 4 (Improper AUE of Clifford + T Circuits; via Theorem 7.1 in [chen2024stabilizer]).
@@
Corollary 5 (Improper AUE of Product Circuits; via Corollary 8.3 in [chen2024stabilizer]).
[…]
@@
Corollary 6 (Improper AUE of Clifford Product Circuits; via Corollary 9.3 in [chen2024stabilizer]).
@@ 4.3 Pauli strings
1:Prepare NN copies of $|v(U)\rangle$ and measure them in the $\{|v(\sigma_{x})\rangle\}_{x}$ basis to get strings $\{x_{i}\}_{i=1}^{N}$
@@
$\left|\hat{p}(x)-|\hat{U}_{x}|^{2}\right|\leq\epsilon_{1},\qquad\forall x\in\mathcal{P}_{n}$ (4.59)
@@
$|\hat{U}_{z}|^{2}+\epsilon_{1}\geq\hat{p}(z)\qquad\text{ and }\qquad\hat{p}(x^{\star})\geq|\hat{U}_{x^{\star}}|^{2}-\epsilon_{1}.$ (4.61)
@@ 4.4 Unitary $k$-juntas
Now, we consider agnostic learning with respect to the class of $n$-qubit unitary $k$-juntas $\mathcal{J}_{k,n}$ [montanaro2008quantum]. An $n$-qubit $k$-junta is a $2^{n}$-dimensional unitary that acts nontrivially on $k$ qubits. More formally, we have the following definition.
@@
The algorithm is reminiscent of other junta learning algorithms [chen2023testing, atici2007quantum] in structure. Namely, we first identify the qubits $S$ that the $k$-junta acts nontrivially on and then learn the unitary on the set $S$. In contrast to [chen2023testing], we use the Pauli weight (Definition 6) rather than the influence to measure how nontrivially a unitary acts on a given qubit. In the realizable setting, both the weight and influence of a subset are equivalent. However, in the agnostic setting, these notions are quite distinct, as demonstrated by the following example.
@@
$\mathcal{P}_{S}=\{x\in\{0,1,2,3\}^{n}\mathrel{\mathop{\ordinarycolon}}\mathsf{supp}(x)\subseteq S\}$ (4.68)
for a subset $S\subseteq[n]$. The resulting learned operator $V_{1}=\sum_{x\in\mathcal{P}_{S}}\alpha_{x}\sigma_{x}$ may not be unitary, so in order to achieve proper agnostic unitary estimation, we need an additional projection which approximates $V_{1}$ by a unitary. For this projection, we use Lemma 14, which takes influence from [huang2024learning]. This algorithm achieves the following guarantee.
@@
$|\hat{w}_{S_{j}}-w_{S_{j}}(U)|\leq\epsilon\qquad\forall j\in[M]$ (4.71)
@@
$\hat{w}_{S}\triangleq\frac{1}{N}\sum_{i\in[N]}b_{S,i},\qquad b_{S,i}\triangleq\mathbf{1}[x_{i}\in\mathcal{P}_{S}].$ (4.72)
@@
[…]
@@
$|\hat{w}_{S_{j}}-w_{S_{j}}(U)|\leq\epsilon\qquad\forall j\in[M]$ (4.78)
@@
with probability at least $1-\delta/2$, for all subsets $S\subseteq[n]$ of cardinality $|S|=k$. By Lemma 15, this requires
@@
In the next step of Algorithm 2, we estimate all $4^{k}$ Pauli coefficients for all Pauli strings supported on $S^{\prime}$. We use Lemma 11 $4^{k}$ times, to obtain estimates $\alpha_{x}$ such that with probability at least $1-\delta/2$,
$|\alpha_{x}-\hat{U}_{x}|\leq\epsilon_{2},\qquad\forall x\in\mathcal{P}_{S^{\prime}}$ (4.91)
@@
[…]